OpenStudy (anonymous):

Rewrite with only sin x and cos x. sin 3x - cos x

OpenStudy (anonymous):

sin(A+B)=sin A cos B + cos A sin B

OpenStudy (anonymous):

These are my options: 2 sin x - sin3x - cos x 2 sin x cos2x + sin x - 2 sin3x - cos x 2 sin3x cos4x + 1 3 sin x cos x - sin3x - cos x

OpenStudy (anonymous):

my method would be breaking 3x in to 2x + x

OpenStudy (anonymous):

Actually my options are: 2 sin x - sin^3x - cos x 2 sin x cos^2x + sin x - 2 sin^3x - cos x 2 sin3x cos^4x + 1 3 sin x cos x - sin^3x - cos x

OpenStudy (cwrw238):

yea - i got that wrong - not my day today

OpenStudy (anonymous):

Can someone actually explain then? Please and thank you :)

OpenStudy (cwrw238):

i'll think i'll come back tomorrow lol

OpenStudy (anonymous):

sin 3x - cos x for now, im going to ignore the - cos x

OpenStudy (anonymous):

sin 3x = sin ( 2x +x) sin(A+B)=sin A cos B + cos A sin B in this case, A = 2x , B = x

OpenStudy (anonymous):

sin ( 2x + x) = sin 2x cos x + cos 2x sin x

OpenStudy (anonymous):

then double angle identities

OpenStudy (anonymous):

That's what I have but then what do we do with the - cos x

OpenStudy (anonymous):

Oh okay

OpenStudy (anonymous):

throw the -cos x onto the entire thing?

OpenStudy (anonymous):

Do we use the sine double angle identity?

OpenStudy (anonymous):

yes because we have a sin 2x we also have a cos 2x that we need to get rid of so we need to use the cosine double angle identity

OpenStudy (whpalmer4):

Use \(\cos 2x = 1-2\sin^2x\) and it comes out nicely from where you've started...

OpenStudy (anonymous):

So it would come out to 2 sin^3x cos^4x + 1?

OpenStudy (anonymous):

@whpalmer4

OpenStudy (anonymous):

@completeidiot

OpenStudy (whpalmer4):

\[\sin 3x - \cos x = \sin(2x+x) -\cos x= \sin 2x\cos x + \cos 2x\sin x - \cos x\]\[\sin 2x\cos x + \cos 2x\sin x - \cos x= (2\sin x \cos x)cos x+\cos 2x\sin x - \cos x\]\[=(2\sin x \cos x)\cos x+(1-2\sin^2x)\sin x - \cos x\]\[= 2\sin x\cos^2x + \sin x - 2\sin^3x - \cos x\]

OpenStudy (whpalmer4):

I like that the \(\cos 2x\) identity gives you 3 choices, so you can head in 3 different directions if needed. I dislike that the \(\cos 2x\) identity gives 3 you choices, because you need to know which direction to go in :-)

OpenStudy (whpalmer4):

(I tried the other 2 first!)